OpenStudy (anonymous):

At the farmer's market, there is a scale that measures the weight of fruit and vegetables. The needle is 16 cm long, and it rotated 68° when Danielle placed six oranges on it. How far did the needle travel as it rotated?

OpenStudy (mathstudent55):

The length of an arc of a circle, s, is \(\Large s = \dfrac{n}{360^o} \times 2 \pi r\) where n = the central angle of the arc, and r = radius

OpenStudy (anonymous):

How do u solve for anser

OpenStudy (mathstudent55):

\(\Large s = \dfrac{n}{360^o} \times 2 \pi r\) \(\Large s = \dfrac{68^o}{360^o} \times 2 \pi (16~cm)\) Just do the math.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

i have four chocies

OpenStudy (lynfran):

let pi=3.142

OpenStudy (mathstudent55):

Do you have a calculator?

OpenStudy (anonymous):